OpenStudy (anonymous):

MEDAL!! Evaluate the expressions: A) arctan(-sqrt3) B) Arctan(-sqrt3)

OpenStudy (anonymous):

hey! I know the answer to A) -pi/3, but I don't know B)

OpenStudy (ikram002p):

well guess the angle wich its tan =-sqrt3

OpenStudy (anonymous):

will B) be different than A)? It is just Arctan instead of arctan. My teacher said there is a subtle difference

hartnn (hartnn):

I don't think there is ANY difference between arctan and Arctan.....its like saying 'not' and 'Not' are different....

OpenStudy (anonymous):

my teacher said there is a difference though. I just don't know if it effects the answer

OpenStudy (ikram002p):

the difference btw A and B is that A could be in the first quirtor or the third B could be in the second or the forth 

OpenStudy (anonymous):

so how does that effect the A) -pi/3 answer?

OpenStudy (anonymous):

@ikram002p

hartnn (hartnn):

and where did you learn that ? i never came across such thing.

OpenStudy (anonymous):

I never actually learned it, but my teacher said there is a difference. http://openstudy.com/study#/updates/4f6b82ace4b014cf77c8494b

OpenStudy (anonymous):

@hartnn do you not know?

hartnn (hartnn):

such notations and not used everywhere, where i learn, arctan and Arctan are certainly the same.

OpenStudy (anonymous):

it seems odd though that the teacher would include both of these equations back to back on a study guide. Surely, both answers aren't the same

OpenStudy (ikram002p):

mmm @washcaps there is no difference btw arc and Arc but angels A ,b are different

OpenStudy (anonymous):

so both answers are -pi/3?

OpenStudy (ikram002p):

one sec there is a diffrence btw arc and Arc sry i dint note that before arc could be any angle but Arc should be an angle with the range [-pi,pi] @hartnn

OpenStudy (anonymous):

oh ok, so how would that effect the answer

OpenStudy (ikram002p):

well arc = -pi/3 or = -pi/3 +2pi or = -pi/ 3 + 2kpi but Arc would only be -pi /3

OpenStudy (anonymous):

so I should write -pi/3 +2kpi for part A to clarify to the teacher the difference

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

yess ur correct !

OpenStudy (anonymous):

ahh thank you so much

OpenStudy (ikram002p):

nope ur wlc anytime !